Crane Company purchased a machine on July 1, 2017, for $900000. The machine has an estimated useful life of five years and a salvage value of $190000. The machine is being depreciated from the date of acquisition by the 150% declining-balance method. For the year ended December 31, 2017, Crane should record depreciation expense on this machine of

$270000.
$180000.
$135000.
$106500.
Answer #1

Answer: $135000.

Depreciation Expense = Cost x 150% x 1/life of asset

For Crane Company = 900000 x 150% x 1/5 = 270000 annually

Till Dec 31 = 270000 / 2 = 135000